OpenStudy (anonymous):

A data consists of the following data points: (3,5), (5,8), (6,13) The slope of the best-fit line is 2.5. Find the y-intercept of this line.

OpenStudy (anonymous):

\(\Large{b_0=\bar{y}-b_1\bar{x}}\)

OpenStudy (anonymous):

\(\Large{b_1}\) is the slope

OpenStudy (anonymous):

So would it be -2?

OpenStudy (anonymous):

to find the y in tercept you do [y1-y2divslope(x1-x2)\] and y1 and x1 stays variblaes with subtitail 1.

OpenStudy (anonymous):

\(\bar{x} =\Large{ \frac{14}{3}}\) and \(\bar{y} =\Large{ \frac{26}{3}}\) so plug in the numbers and compute. i don't get -2...

OpenStudy (anonymous):

to which one do i plug the number in?

OpenStudy (anonymous):

\(b_0\) is the y intercept. the formula is:\[b_0=\bar{y}-b_1 \bar{x}\] you have eveything, you just need to compute.

OpenStudy (anonymous):

are you in a stats class?

OpenStudy (anonymous):

No, it's Algebra 1.

OpenStudy (anonymous):

really? did they give a formula for this?

OpenStudy (anonymous):

No, that's why I'm really confused.

OpenStudy (anonymous):

understandably! well, now you have the formula... you just have to plug in the numbers and compute.

OpenStudy (anonymous):

\[b_0=\frac{ 26 }{ 3 }-\frac{ 5 }{ 2 }\cdot \frac{ 14 }{ 3 }\]

OpenStudy (anonymous):

I got 504/9